Use factor theorem to solve 2x^3+7x^2+2x-3

Answer by ewatrrr(24785)


Hi,
Note: x = -1 is a solution (-2 + 7 -2 - 3 = 0)
2x^2 + 5x -3
_______________________
x +1 | 2x^3+7x^2+2x-3
2x^2 + 5x -3
(2x - 1)(x + 3 )= 0
(2x - 1)=0
x = 1/2
(x + 3 )= 0
x = -3
solution set is {-3, -1, 1/2}